### v2

# The geometric control of boundary-catalytic branching processes

Abstract
Boundary-catalytic branching processes describe a broad class of natural phenomena where the population of diffusing particles grows due to their spontaneous binary branching (e.g., division, fission or splitting) on a catalytic boundary located in a complex environment. We investigate the possibility of the geometric control of the population growth by compensating the proliferation of particles due to catalytic branching events by their absorptions in the bulk or on absorbing regions of the boundary. We identify an appropriate Steklov spectral problem to obtain the phase diagram of this out-of-equilibrium stochastic process. The principal eigenvalue determines the critical line that separates an exponential growth of the population from its extinction in a bounded domain. In other words, we establish a powerful tool for calculating the optimal absorption rate that equilibrates the opposite effects of branching and absorption events and thus results in steady-state behavior of this diffusion-reaction system. Moreover, we show the existence of a critical catalytic rate above which no compensation is possible, so that the population cannot be controlled and keeps growing exponentially. The proposed framework opens promising perspectives for better understanding, modeling and control of various boundary-catalytic branching processes, with applications in physics, chemistry, and life sciences.
branching processes, diffusion-mediated phenomena, catalysis, biochemical reactions, boundary local time, Steklov problem
pacs:
02.50.-r, 05.40.-a, 02.70.Rr, 05.10.Gg
I Introduction
Branching processes are critically important in disciplines as diverse as atomic energy (neutron production in a nuclear reactor), optics (stimulated emission in a laser cavity), chemistry (heterogeneous catalysis), biology (bacterial colony growth), ecology (population dynamics), social sciences (genealogy), to name but a few [1, 2, 3]. In the simplest homogeneous setting, each particle can spontaneously split into two identical copies that undergo the same branching process independently of each other. These stochastic branching events result in an exponential growth of the population size in time. In most applications in Nature and industry, however, branching processes are spatially heterogeneous and occur on catalytic sites [4, 5, 6, 7]. Such catalytic branching processes strongly rely on the motion of particles and thus couple branching mechanisms to particle dynamics. For instance, when a bacterial colony explores the available space foraging for food, the population growth is controlled by available resources, their amount and spatial configuration.
In this paper, we consider the so-called boundary-catalytic branching (BCB) processes when particles diffuse in a restricted environment and branching events occur at boundaries: physical, spatial, or functional. This boundary-driven behavior appears in many real-world settings. In physics, such models describe particle systems where reactions occur only at interfaces, such as catalytic surfaces in chemical kinetics [8, 9, 10, 11]. In life sciences, they help explaining phenomena like cell proliferation at tissue boundaries, stem-cell niche dynamics, the spread of populations or epidemics when reproduction or contamination is concentrated at ecological edges [12, 13, 14]. By combining stochastic branching with spatial constraints, boundary-catalytic models reveal how local interactions at borders can shape global system behavior, making them valuable tools for understanding complex, heterogeneous environments across scientific disciplines.
Inspired by applications in life sciences, we investigate a critically relevant possibility of the geometric control of the population size: how can one incorporate absorbing regions into the environment to efficiently eliminate excessive particles in order to put an explosive population growth under control? For this purpose, we employ various tools such as a probabilistic construction of the BCB process based on the boundary local time, a spectral representation of the mean population size and its long-time asymptotic behavior, matched asymptotic expansions, and optimization schemes for choosing the appropriate reactivity to balance branching and absorption events. Despite intensive research on diffusion-mediated phenomena and related first-passage statistics [15, 16, 17, 18, 19, 20, 21], the important topic of BCB processes is yet unexplored in physics literature. In fact, even though the general concept of balancing birth and death events is very natural and has been studied in the past (see, e.g., [22, 23, 24] for the analysis of phase diagrams for branching-annihilating random walks), the spatial location of branching and absorption events on a boundary brings a new dimension to this fundamental problem.
II Model
We consider the following model of BCB processes. At time 0, a single particle is released from a point $\bm{x}$ and diffuses inside a confining domain $\Omega\subset{\mathbb{R}}^{d}$ with a constant diffusivity $D$. The boundary of this environment, $\partial\Omega=\Gamma_{\mathrm{c}}\cup\Gamma_{\mathrm{a}}\cup\Gamma_{\mathrm{r}}$, is in general partitioned into three disjoint subsets: a catalytic region $\Gamma_{\mathrm{c}}$, an absorbing region $\Gamma_{\mathrm{a}}$, and the remaining reflecting region $\Gamma_{\mathrm{r}}$. As illustrated on Fig. 1(a), both catalytic and absorbing regions can be composed of multiple “pieces”, either located on the boundary or in the bulk. The absorbing region is characterized by a reactivity $\kappa_{\mathrm{a}}=q_{\mathrm{a}}D\geq 0$, which can range from 0 (no absorption) to $\infty$ (a perfect sink that instantly kills the particle upon its first arrival onto $\Gamma_{\mathrm{a}}$). When $q_{\mathrm{a}}$ is finite, each particle arrived onto the absorbing region $\Gamma_{\mathrm{a}}$ is either absorbed (with a small probability proportional to $q_{\mathrm{a}}$), or reflected back to resume its diffusion in $\Omega$ [25, 26, 27, 28]. As a result, the particle is killed when the random number of its encounters with the absorbing region, represented by the boundary local time $\ell_{t}^{\mathrm{a}}$ on $\Gamma_{\mathrm{a}}$, exceeds an independent random threshold $\hat{\ell}^{\mathrm{a}}$ obeying the exponential distribution ${\mathbb{P}}\{\hat{\ell}^{\mathrm{a}}>\ell\}=e^{-q_{\mathrm{a}}\ell}$ [29]. The parameter $q_{\mathrm{a}}$ quantifies thus the intensity of absorption events. With some abuse of language, we will call $q_{\mathrm{a}}$ the absorption rate, even though $q_{\mathrm{a}}$ has units of the inverse of length, in accordance to the fact that the boundary local time $\ell_{t}^{\mathrm{a}}$ has units of length. A similar construction was employed in the probabilistic description of permeation processes such as snapping out Brownian motion [30, 31, 32].
In the same vein, the catalytic region $\Gamma_{\mathrm{c}}$ is characterized by a reactivity $\kappa_{\mathrm{c}}=q_{\mathrm{c}}D\geq 0$, so that the particle hitting $\Gamma_{\mathrm{c}}$ can either be replaced by two new particles (with a small probability proportional to $q_{\mathrm{c}}$), or be reflected back to resume its diffusion (Fig. 1(b)). In other words, the branching event is triggered when the boundary local time $\ell_{t}^{\mathrm{c}}$ of the particle on $\Gamma_{\mathrm{c}}$ exceeds another independent random threshold $\hat{\ell}^{\mathrm{c}}$ obeying the exponential distribution ${\mathbb{P}}\{\hat{\ell}^{\mathrm{c}}>\ell\}=e^{-q_{\mathrm{c}}\ell}$, with the catalytic rate (or branching intensity) $q_{\mathrm{c}}$. The two newborn particles are released from the position of their branching event (with their boundary local times set to 0) and diffuse independently of each other, until the next branching or absorption event, and so on (see [33, 34, 35] for further mathematical details). The competition between the opposite effects of catalytic and absorbing regions will determine the population dynamics and, in particular, the mean population size $N(t|\bm{x})$ (i.e., the average number of particles at time $t$). Even though the above probabilistic constructions of absorption and branching events are identical, we will show that the impact of the rates $q_{\mathrm{a}}$ and $q_{\mathrm{c}}$ onto the population size is dramatically different.
Figure 1: (a) Schematic view of a confining domain $\Omega$, whose boundary $\partial\Omega$ is split into three disjoint sets: catalytic region $\Gamma_{\mathrm{c}}$, absorbing region $\Gamma_{\mathrm{a}}$, and reflecting region $\Gamma_{\mathrm{r}}$, each of them may be composed of a finite number of disjoint subsets (e.g., $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{a}}^{1}\cup\Gamma_{\mathrm{a}}^{2}\cup\Gamma_{\mathrm{a}}^{3}$ is shown). As the boundary $\partial\Omega$ does not need to be connected, both $\Gamma_{\mathrm{c}}$ and $\Gamma_{\mathrm{a}}$ can be either patches on the outer boundary (as shown for $\Gamma_{\mathrm{a}}$), or surfaces of impenetrable solids in the bulk (as shown for $\Gamma_{\mathrm{c}}$), or their combinations. (b) A random realization of the BCB process on the interval $\Omega=(0,L)$ with $\Gamma_{\mathrm{c}}=\{0\}$ and $\Gamma_{\mathrm{a}}=\{1\}$, $\kappa_{\mathrm{a}}=1$, $\kappa_{\mathrm{c}}=1$, $D=1$, and $x=0.5$ (here $\Gamma_{\mathrm{r}}=\emptyset$). Two branching events are indicated by half-circles, one absorption event is indicated by a cross.
III Results
III.1 Negative reactivity on the catalytic region
As the number of diffusing particles and thus the dimension of the configuration space randomly change in time, a probabilistic description of branching processes requires elaborate mathematical tools such as measure-valued stochastic processes (or superprocesses) and nonlinear PDEs [36, 37, 38, 39, 40]. In turn, the mean population size $N(t|\bm{x})$ still admits the linear PDE description via the diffusion equation with Robin boundary condition,
$\partial_{t}N(t|\bm{x})$ $=D\Delta N(t|\bm{x})\quad\textrm{in}~\Omega,$ (1a)
$-\partial_{n}N(t|\bm{x})$ $=q(\bm{x})N(t|\bm{x})\quad\textrm{on}~\partial\Omega,$ (1b)
with the initial condition $N(0|\bm{x})=1$. Here $\partial_{n}$ is the normal derivative oriented outward the domain $\Omega$, $\Delta$ is the Laplace operator governing the diffusion dynamics, and $q(\bm{x})$ is the piecewise-constant function: $q(\bm{x})=q_{\mathrm{a}}$ on the absorbing region $\Gamma_{\mathrm{a}}$, $q(\bm{x})=0$ on the reflecting region $\Gamma_{\mathrm{r}}$, and $q(\bm{x})=-q_{\mathrm{c}}$ on the catalytic region $\Gamma_{\mathrm{c}}$. The negative parameter $q_{\mathrm{c}}$ in the Robin boundary condition incorporates the catalytic effect of $\Gamma_{\mathrm{c}}$, which plays the role of a source of particles that are released into the bulk after branching. In fact, the left-hand side of Eq. (1b) is the net diffusive flux to the bulk, which is set to be equal to the flux of created particles, the latter being proportional to $N(t|\bm{x})$. The PDE description (1) in the purely catalytic case (i.e., $\partial\Omega=\Gamma_{\mathrm{c}}$) was rigorously established in [33] for polygonal bounded domains but its extension to our mixed setting with a piecewise constant rate $q(\bm{x})$, other domains and higher dimensions seems to be straightforward. Moreover, as branching and absorption events occur exclusively on the boundary, we expect that ordinary diffusion can be replaced by a general Markov process governed by a second-order elliptic operator, in which case the diffusion equation (1a) has to be substituted by the backward Fokker-Planck (or Kolmogorov) equation. A mathematical justification of such an extension and the analysis of its consequences are beyond the scope of this paper.
The presence of the negative parameter $q_{\mathrm{c}}$ makes a crucial difference from the conventional analysis of the survival probability that admits the same PDE description and characterizes first-passage times, target search problems, and diffusion-controlled reactions in reactive media [15, 16, 12, 13, 17, 9, 10, 18, 19, 14, 20, 21, 11]. In this study, we aim at answering how the geometric configuration of absorbing and catalytic regions controls the long-time behavior of the mean population size $N(t|\bm{x})$. We will treat separately the cases of bounded and unbounded environments.
III.2 The geometric control of the population growth
The diffusive dynamics is governed by the Laplace operator and its spectral properties. When the domain $\Omega$ is bounded, the spectrum is discrete, i.e., there are infinitely many eigenpairs $\{\lambda,u\}$, satisfying
$-\Delta u$ $=\lambda u~~\textrm{in}~\Omega,\qquad\partial_{n}u-q_{\mathrm{c}}u=0~~\textrm{on}~\Gamma_{\mathrm{c}},$ (2a)
$\partial_{n}u+q_{\mathrm{a}}u$ $=0~~\textrm{on}~\Gamma_{\mathrm{a}},\qquad\partial_{n}u=0~~\textrm{on}~\Gamma_{\mathrm{r}}.$ (2b)
The eigenvalues are enumerated by the index $k=0,1,2,\cdots$ to form an increasing sequence, $\lambda_{0}\leq\lambda_{1}\leq\cdots\nearrow+\infty$, whereas the eigenfunctions $\{u_{k}\}$ form a complete orthonormal basis of $L^{2}(\Omega)$ [41]. As a consequence, the solution of the diffusion equation (1) admits a standard spectral expansion,
$N(t|\bm{x})=\sum\limits_{k=0}^{\infty}e^{-Dt\lambda_{k}}\,u_{k}(\bm{x})\int\limits_{\Omega}u_{k}.$ (3)
It is therefore clear that the long-time behavior of $N(t|\bm{x})$ is controlled by the principal (smallest) eigenvalue $\lambda_{0}$: $N(t|\bm{x})\propto e^{-Dt\lambda_{0}}$ as $t\to\infty$. In sharp contrast to the survival probability, which always decays exponentially due to eventual absorption on $\Gamma_{\mathrm{a}}$ with $\kappa_{\mathrm{a}}>0$, the behavior of the mean population size depends on the competition between the proliferative effect of branching events on $\Gamma_{\mathrm{c}}$ and the destroying effect of absorption events on $\Gamma_{\mathrm{a}}$. In other words, $N(t|\bm{x})$ may exponentially decay ($\lambda_{0}>0$), reach a steady-state limit ($\lambda_{0}=0$), or exponentially grow ($\lambda_{0}<0$). The geometric control of the population growth has therefore two aspects: (i) for a given geometric setting, to find the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},\lambda_{0})$ that yields the desired value of $\lambda_{0}$ for a given catalytic rate $q_{\mathrm{c}}$; (ii) for given $q_{\mathrm{a}}$ and $q_{\mathrm{c}}$, to find the optimal geometric configuration of absorption regions to achieve the desired $\lambda_{0}$. In the following, we will mainly focus on the first aspect for the target value $\lambda_{0}=0$ ensuring the steady-state limit and thus separating the exponential growth from extinction on the phase diagram in the $(q_{\mathrm{c}},q_{\mathrm{a}})$ space.
In the low-rate regime when both $q_{\mathrm{c}}$ and $q_{\mathrm{a}}$ are small, Robin boundary conditions in Eqs. (2) are close to the Neumann one so that the principal eigenvalue $\lambda_{0}$ is close to zero, whereas the associated eigenfunction $u_{0}$ is close to a constant. A standard perturbation approach yields in the leading order (see Appendix A for details)
$\lambda_{0}\approx\frac{q_{\mathrm{a}}|\Gamma_{\mathrm{a}}|-q_{\mathrm{c}}|\Gamma_{\mathrm{c}}|}{|\Omega|}\,$, (4)
where $|\Omega|$ is the volume of the confining domain $\Omega$, while $|\Gamma_{\mathrm{a}}|$ and $|\Gamma_{\mathrm{c}}|$ are the surface areas of the absorbing and catalytic regions, respectively. In this regime, the production and the destruction of particles are proportional to the surface areas of the respective regions, and $\lambda_{0}$ quantifies the competition between these two opposite mechanisms. In order to balance branching and absorption events and to reach the steady-state behavior with $\lambda_{0}=0$, the simple relation (4) suggests to fix the optimal absorption rate as
$\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)\approx q_{\mathrm{c}}\frac{|\Gamma_{\mathrm{c}}|}{|\Gamma_{\mathrm{a}}|}\,.$ (5)
In the following, we investigate the geometric control beyond this approximate relation and reveal its limitations.
III.3 Explicit examples
To gain some intuition onto the geometric control, let us first inspect the case of an interval of length $L$, $\Omega=(0,L)$, with $\Gamma_{\mathrm{c}}=\{0\}$, $\Gamma_{\mathrm{a}}=\{L\}$ and $\Gamma_{\mathrm{r}}=\emptyset$ (no reflecting region). The Laplacian eigenfunctions are obtained as linear combinations of sine and cosine functions in a standard way (see Appendix B for details); in particular, the eigenvalues $\lambda_{k}$ are the solutions of the transcendental equation:
$(\lambda_{k}+q_{\mathrm{a}}q_{\mathrm{c}})\frac{\tan(\sqrt{\lambda_{k}}L)}{\sqrt{\lambda_{k}}}=q_{\mathrm{a}}-q_{\mathrm{c}}.$ (6)
When $q_{\mathrm{c}}=0$ (no branching, just reflections at $x=0$), one retrieves the classical problem of the survival of a diffusing particle on the interval with one absorbing and one reflecting endpoint [42]. In turn, if $q_{\mathrm{c}}=q_{\mathrm{a}}$, Eq. (6) has infinitely many positive solutions $\lambda_{k}=\pi^{2}k^{2}/L^{2}$, with $k=1,2,\cdots$, as well as one negative solution: $\lambda_{0}=-q_{\mathrm{a}}q_{\mathrm{c}}$. One sees that, even though branching and absorption events are governed by the same mechanism, the catalytic effect always “wins” in this setting, yielding $\lambda_{0}<0$. In order to balance the catalytic effect, one therefore needs absorption events at higher rate. Setting $\lambda_{k}=0$ in Eq. (6) yields $q_{\mathrm{a}}q_{\mathrm{c}}L=q_{\mathrm{a}}-q_{\mathrm{c}}$ so that the steady-state regime can only emerge at the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)=q_{\mathrm{c}}/(1-q_{\mathrm{c}}L)$ and only if $q_{\mathrm{c}}\leq q_{\mathrm{c}}^{\rm crit}=1/L$. In fact, if $q_{\mathrm{c}}$ exceeds the critical value $q_{\mathrm{c}}^{\rm crit}$, even the perfect sink with $q_{\mathrm{a}}=\infty$ cannot compensate the catalytic effect, implying the exponential growth of the mean population size.
The above explicit analysis can be extended to higher dimensions $d\geq 2$ by considering diffusion in a spherical shell $\Omega=\{\bm{x}\in{\mathbb{R}}^{d}~:~R<|\bm{x}|<L\}$ between a catalytic sphere of radius $R$ with $q_{\mathrm{c}}>0$ and a concentric absorbing sphere of radius $L$ with $q_{\mathrm{a}}\geq 0$. Solving Eqs. (2) with $\lambda=0$ via the separation of variables, we get (see Appendix C for details):
$\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)=q_{\mathrm{c}}\frac{(R/L)^{d-1}}{1-q_{\mathrm{c}}/q_{\mathrm{c}}^{\rm crit}}\,$, (7)
where
$q_{\mathrm{c}}^{\rm crit}=\frac{1}{R}\times\begin{cases}1/\ln(L/R)\hskip 65.44133pt(d=2),\cr(d-2)/(1-(R/L)^{d-2})\quad(d\geq 3).\end{cases}$ (8)
In the low-rate catalytic regime ($q_{\mathrm{c}}\ll q_{\mathrm{c}}^{\rm crit}$), one retrieves the approximate relation (5), which for a spherical shell gives $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)\approx q_{\mathrm{c}}(R/L)^{d-1}$, as in Eq. (7). However, even though the absorbing region can be very large, there exists the critical value $q_{\mathrm{c}}^{\rm crit}$ of the catalytic rate, above which the catalytic production cannot be compensated, yielding an exponential growth of the population. Curiously, $q_{\mathrm{c}}^{\rm crit}$ as a function of $R$ exhibits a minimum at $R_{0}/L=(d-1)^{-1/(d-2)}$ (and $R_{0}/L=e^{-1}$ in two dimensions), i.e., there is an optimal radius of the catalytic region, at which the branching events are the most proliferative and thus more difficult to compensate.
III.4 Phase diagram for bounded domains
Most importantly, the above intuitive picture remains valid for general bounded domains. The eigenvalue problem (2) that we used to determine the principal eigenvalue $\lambda_{0}$, can also be regarded as a mathematical tool for the geometric control of BCB processes. In fact, fixing the catalytic rate $q_{\mathrm{c}}$ and the desired value of $\lambda$, one can treat $q_{\mathrm{a}}$ as a spectral parameter, denoted as $\sigma$, and search for eigenpairs $\{\sigma,u\}$ satisfying Eqs. (2). This is the so-called generalized Steklov spectral problem, which is known to have a discrete spectrum [41]. Due to the presence of the catalytic region $\Gamma_{\mathrm{c}}$ with the negative Robin parameter $q_{\mathrm{c}}$, the principal (smallest) eigenvalue $\sigma_{0}$ is strictly negative. This eigenvalue determines the optimal absorption rate, $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},\lambda_{0})=-\sigma_{0}$, that one has to use to achieve the desired value $\lambda_{0}$ in a given geometric setting. In the particular case $\lambda_{0}=0$, the function $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ determines the critical line in the $(q_{\mathrm{c}},q_{\mathrm{a}})$ plane that separates the extinction regime ($q_{\mathrm{a}}>\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$) from the growth regime ($q_{\mathrm{a}}<\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$) on the phase diagram. Moreover, the associated eigenfunction $u_{0}(\bm{x})$ determines the steady-state spatial profile of the mean population size $N(\infty|\bm{x})$ (further spectral insights are discussed in Appendix D). Alternatively, one could fix $q_{\mathrm{a}}$ and treat $q_{\mathrm{c}}$ as a spectral parameter $\mu$ to determine the function $\hat{q}_{\mathrm{c}}(q_{\mathrm{a}},\lambda_{0})$ that fixes $q_{\mathrm{c}}$ to achieve the desired value of $\lambda_{0}$ for a given $q_{\mathrm{a}}$. For instance, if we are interested in the steady-state solution ($\lambda_{0}=0$) with the maximal absorption rate $q_{\mathrm{a}}=\infty$ (a perfect sink), Eq. (2) is reduced to
$\Delta v$ $=0\quad\textrm{in}~\Omega,\qquad\partial_{n}v=\mu v\quad\textrm{on}~\Gamma_{\mathrm{c}},$ (9a)
$\partial_{n}v$ $=0\quad\textrm{on}~\Gamma_{\mathrm{r}},\qquad v=0\quad\textrm{on}~\Gamma_{\mathrm{a}},$ (9b)
and the principal eigenvalue $\mu_{0}$ gives the critical catalytic rate $q_{\mathrm{c}}^{\rm crit}$ (here the eigenpairs are denoted as $\{\mu,v\}$ to distinguish them from the previous problem). Moreover, we show (see Appendix E) that the associated eigenfunction $v_{0}$ determines the asymptotic behavior of $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ as $q_{\mathrm{c}}$ approaches $q_{\mathrm{c}}^{\rm crit}$:
$\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)\approx\frac{Q^{2}}{q_{\mathrm{c}}^{\rm crit}-q_{\mathrm{c}}}\,,\qquad Q=\frac{\|\partial_{n}v_{0}\|_{L^{2}(\Gamma_{\mathrm{a}})}}{\|v_{0}\|_{L^{2}(\Gamma_{\mathrm{c}})}}\,$, (10)
which agrees with Eq. (7) for a spherical shell. The identification of this general mathematical framework for relating the parameters $q_{\mathrm{a}}$, $q_{\mathrm{c}}$ and $\lambda_{0}$ and for determining $q_{\mathrm{c}}^{\rm crit}$ opens promising perspectives for the geometric control using spectral geometry and related asymptotic and numerical tools.
Figure 2: (a,b) Two configurations composed of one catalytic circle of radius $R=0.1$ at the center, and 9 randomly located absorbing circles of radius $R_{\mathrm{a}}$ (with $R_{\mathrm{a}}=0.05$ (a) and $R_{\mathrm{a}}=0.1$ (b)), enclosed by a reflecting circle of radius $L=1$. Colors ranging from dark blue to dark red represent variations of the limiting mean population size $N(\infty|\bm{x})$ for the special case $q_{\mathrm{a}}=\infty$ (perfect sinks) and $q_{\mathrm{c}}=q_{\mathrm{c}}^{\rm crit}$. (c) The optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ as a function of the catalytic rate $q_{\mathrm{c}}$ for two considered configurations (shown by solid line for $R_{\mathrm{a}}=0.05$ and by dashed line for $R_{\mathrm{a}}=0.1$). Dotted lines show the low-rate approximation (5), whereas dash-dotted lines present the asymptotic relation (10), with $Q\approx 15.78$ for $R_{a}=0.05$ and $Q\approx 19.21$ for $R_{a}=0.1$. Vertical lines indicate the critical values $q_{\mathrm{c}}^{\rm crit}\approx 6.61$ and $q_{\mathrm{c}}^{\rm crit}\approx 7.62$ for two configurations. These numerical results were obtained by a finite-element method [43], with the maximal meshsize $h_{\rm max}=0.01$ (see Appendix F for details).
Figure 2(c) shows the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ for two configurations (illustrated on panels 2(a,b)) with one catalytic region and nine randomly located absorbing regions of circular shape. This rate was computed numerically by solving the Steklov spectral problem via a finite-element method [43] (see Appendix F). In the low-rate regime, the approximate relation (5) is valid and provides the lower bound on $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$, as expected. In turn, as $q_{\mathrm{c}}$ approaches the critical value $q_{\mathrm{c}}^{\rm crit}$ (shown by vertical dashed lines), $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ diverges according to Eq. (10). As said earlier, Fig. 2(c) can be interpreted as the phase diagram in the space of $q_{\mathrm{c}}$ and $q_{\mathrm{a}}$ rates, in which the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ is the critical line that separates an exponential growth of population from its extinction.
III.5 Asymptotic results for small regions
When the catalytic and absorbing regions are small (Fig. 2(a,b)), matched asymptotic expansions can be applied to estimate the principal Steklov eigenvalue $\mu_{0}$ and thus the critical value $q_{\mathrm{c}}^{\rm crit}$ without solving the spectral problem (9). In two dimensions, the asymptotic analysis of the Steklov problem (9) was undertaken in [44] for the case when the reflecting boundary $\partial\Omega$ is smooth, connected and covered by small catalytic and absorbing regions. If the catalytic region is a single arc-like set, whereas the absorbing region is a finite union of well-separated arc-like sets, the critical catalytic rate is
$q_{\mathrm{c}}^{\rm crit}=\mu_{0}\simeq\frac{1}{\epsilon_{\rm c}\bigl[C\ln(L/\epsilon_{\rm c})+C_{0}\bigr]}\,$, (11)
where $2L$ is the diameter of the domain $\Omega$, $2\epsilon_{\rm c}$ is the perimeter of the catalytic region $\Gamma_{\rm c}$, $C_{0}=\frac{2}{\pi}\bigl(\frac{3}{2}-\ln 2\bigr)$ is a numerical constant, whereas $C$ depends on the lengths of all patches, their locations on the boundary $\partial\Omega$, and the shape of the confining domain (see Eq. (35) and Appendix G for details). For instance, if the domain $\Omega$ is the disk and the absorbing region is an arc of length $2\epsilon_{\rm a}$, the factor $C$ can be found explicitly, yielding
$q_{\mathrm{c}}^{\rm crit}\approx\frac{\pi}{2\epsilon_{\rm c}}\biggl(-\ln(\epsilon_{\rm c}\epsilon_{\rm a})+\frac{3}{2}+2\ln|\bm{x}_{\rm c}-\bm{x}_{\rm a}|\biggr)^{-1},$ (12)
where $\bm{x}_{\rm c}$ and $\bm{x}_{\rm a}$ are the positions of the catalytic and absorbing regions on the boundary. Expectedly, when the catalytic region $\Gamma_{\mathrm{c}}$ is reduced (i.e., $\epsilon_{\rm c}\to 0$), the critical catalytic rate $q_{\mathrm{c}}^{\rm crit}$ increases, so that it is easier to compensate the effect of branching events. In contrast, if the absorbing region $\Gamma_{\mathrm{a}}$ is reduced (i.e., $\epsilon_{\rm a}\to 0$), less particles can be trapped, and $q_{\mathrm{c}}^{\rm crit}$ decreases. Similarly, the critical catalytic rate also decreases when the distance $|\bm{x}_{\mathrm{c}}-\bm{x}_{\mathrm{a}}|$ between two regions increases. However, both effects are weak due to the logarithmic nature of the terms $\ln(1/\epsilon_{\rm a})$ and $\ln|\bm{x}_{\mathrm{c}}-\bm{x}_{\mathrm{a}}|$. In Appendix G, we provide an extension of this asymptotic analysis to the case when the catalytic region is a small disk inside $\Omega$, whereas the absorbing region is a finite union of small well-separated disks, like the configurations shown in Fig. 2(a,b). A similar asymptotic analysis of the Steklov problem (9) in three dimensions is provided [45].
III.6 Extension to unbounded domains
Up to this point, we focused on bounded domains. The explicit description (7, 8) of the steady-state solution in a spherical shell helps to clarify what happens for unbounded domains when the absorbing region $\Gamma_{\mathrm{a}}$ is (re)moved to infinity. Looking at the limit of Eq. (8) as $L\to\infty$, one can appreciate the crucial difference between two-dimensional and higher-dimensional settings. In two dimensions, the critical value $q_{\mathrm{c}}^{\rm crit}$ logarithmically vanishes as $L\to\infty$. In other words, if the absorbing boundary $\Gamma_{\mathrm{a}}$ is pushed to infinity, one gets $q_{\mathrm{c}}^{\rm crit}=0$ so that, for any $q_{\mathrm{c}}>0$, the population grows exponentially. This is a consequence of the recurrent nature of Brownian motion in the plane: each particle never stops returning to the catalytic surface that surely triggers its branching. The situation is different in higher dimensions, where $q_{\mathrm{c}}^{\rm crit}\to(d-2)/R$ as $L\to\infty$. Even though the absorbing region is removed, eventual escapes of particles to infinity due to the transient character of Brownian motion substitute the absorption events. According to the explicit computation of $N(t|\bm{x})$ for the exterior of a three-dimensional ball of radius $R$ (see Appendix H for details), one can distinguish three scenarios: (i) if $q_{\mathrm{c}}>1/R$, the branching events are more efficient than the escape events, so that the mean population size $N(t|\bm{x})$ grows exponentially, despite the unboundedness of the environment; (ii) if $q_{\mathrm{c}}<1/R$, the mean population size reaches a finite steady-state value $N(\infty|\bm{x})=1+(R/|\bm{x}|)/(1/(q_{\mathrm{c}}R)-1)$; (iii) in the critical setting $q_{\mathrm{c}}=1/R$, the mean population size exhibits an unlimited power-law growth: $N(t|\bm{x})\propto\sqrt{Dt}/|\bm{x}|$, as $t\to\infty$. Moreover, the critical value $1/R$ can be identified again with the principal eigenvalue $\mu_{0}$ of the Steklov problem (9) in the exterior of a ball, for which $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{r}}=\emptyset$.
While the above picture was based on the explicit solution for the exterior of a ball, it remains valid for general exterior domains. In fact, let $\mu_{0}$ be the principal Steklov eigenvalue in the exterior $\Omega={\mathbb{R}}^{d}\backslash\overline{\Omega_{0}}$ of a bounded domain $\Omega_{0}$ with boundary $\partial\Omega=\Gamma_{\mathrm{c}}$ (see [46, 47] for its rigorous formulation). The Laplace operator in $\Omega$ with Robin boundary condition on $\Gamma_{\mathrm{c}}$ has (at least) one strictly negative eigenvalue, if and only if $q_{\mathrm{c}}>q_{\mathrm{c}}^{\rm crit}=\mu_{0}$ [48]. As a consequence, even though the spectral expansion (3) is not applicable because the Laplace operator has the essential spectrum $0,+\infty)$, the presence of the isolated eigenvalue $\lambda_{0}<0$ ensures an exponential growth of the mean population size with the rate $1/(D|\lambda_{0}|)$ (see [49, 50, 51] for further discussions of the related spectral properties). In contrast, the analysis of the long-time asymptotic behavior of the mean population size in the case $q_{c}\leq\mu_{0}$ remains an open problem.
Throughout this paper, we assumed that the particles could be killed exclusively on the absorbing region $\Gamma_{\mathrm{a}}$. In many applications, the bulk medium itself can be aggressive and lead to eventual destructions of the diffusing particles with some rate $\nu>0$ [52, 53, 54]. This mechanism, which can also account for a finite random lifetime of particles and their spontaneous death in the bulk, can be easily incorporated by adding the reactive term $\nu N(t|\bm{x})$ to the right-hand side of the diffusion equation (1). In this case, the steady-state regime would correspond to $\lambda_{0}=-\nu/D$ instead of $\lambda_{0}=0$, and the above analysis can be adapted to this more general situation.
IV Conclusions
In summary, we formulated and solved the fundamental problem of the geometric control of the mean population size of BCB processes. In fact, we identified the generalized Steklov problem as the proper mathematical tool to determine the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},\lambda_{0})$ that ensures the desired behavior of the mean population size at long times. We showed that the absorption events can compensate the branching events to reach a nontrivial steady-state limit only if the catalytic rate $q_{\mathrm{c}}$ does not exceed the critical value $q_{\mathrm{c}}^{\rm crit}$. The latter depends on the geometric configuration of both catalytic and absorbing regions. In the typical case of small regions, the asymptotic formula for $q_{\mathrm{c}}^{\rm crit}$ was derived. These theoretical advances open promising ways for engineering optimal configurations of absorption regions to put the population size under control in various diffusion-reaction systems. Moreover, many other aspects of BCB processes, such as their first-passage time statistics, eventual instabilities due to nonlinear nature of branching events, the rate of relaxation to the steady-state solution (controlled by the next eigenvalue $\lambda_{1}$), role of fluctuations in this critical regime, and extreme value statistics, remain unknown and present open challenging problems for statistical physics and related disciplines.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Acknowledgements.
D.S.G. acknowledges the Simons Foundation for supporting his sabbatical sojourn in 2024 at the CRM (University of Montréal, Canada), as well as the Alexander von Humboldt Foundation for support within a Bessel Prize award.
Appendix A Low-rate regime
In this Section, we provide some details on the perturbation approach in the low-rate regime when both rates $q_{\mathrm{a}}$ and $q_{\mathrm{c}}$ are small. In this setting, the Laplace operator with mixed boundary conditions (2) can be seen as a perturbation of the Laplace operator in the same bounded domain with Neumann boundary condition on the whole boundary $\partial\Omega$. As a consequence, one can search the principal eigenpair $\{\lambda_{0},u_{0}\}$ of the spectral problem (2) as
$\lambda_{0}$ $=\lambda^{(0)}+\epsilon\lambda^{(1)}+\epsilon^{2}\lambda^{(2)}+\cdots,$ (13a)
$u_{0}$ $=u^{(0)}+\epsilon u^{(1)}+\epsilon^{2}u^{(2)}+\cdots,$ (13b)
where $\epsilon$ is a small parameter, and superscripts are used to indicate the correction orders in powers of $\epsilon$. The principal eigenvalue of the “unperturbed” operator is zero, $\lambda^{(0)}=0$, whereas the associated eigenfunction is constant: $u^{(0)}=const$. Substituting the above expansions into Eqs. (2), setting $q_{\mathrm{a}}=a\epsilon$ and $q_{\mathrm{c}}=c\epsilon$, and collecting the terms of the same order, we get in the first order:
$-\Delta u^{(1)}$ $=\lambda^{(1)}\quad\textrm{in}~\Omega,\quad\partial_{n}u^{(1)}=0\quad\textrm{on}~\Gamma_{\mathrm{r}},$ (14a)
$\partial_{n}u^{(1)}$ $=c\quad\textrm{on}~\Gamma_{\mathrm{c}},\quad\partial_{n}u^{(1)}=-a\quad\textrm{on}~\Gamma_{\mathrm{a}}.$ (14b)
Integrating the above Poisson equation over $\Omega$ and applying the Green’s formula, one has
$|\Omega|\lambda^{(1)}=\int\limits_{\Omega}\Delta u^{(1)}=\int\limits_{\partial\Omega}\partial_{n}u^{(1)}=c|\Gamma_{\mathrm{c}}|-a|\Gamma_{\mathrm{a}}|.$
Multiplication of this relation by $\epsilon$ yields Eq. (4), in the first order in $\epsilon$.
In the next order, we have
$-\Delta u^{(2)}$ $=\lambda^{(2)}+\lambda^{(1)}u^{(1)}\quad\textrm{in}~\Omega,\quad\partial_{n}u^{(2)}=0\quad\textrm{on}~\Gamma_{\mathrm{r}},$
$\partial_{n}u^{(2)}$ $=cu^{(1)}\quad\textrm{on}~\Gamma_{\mathrm{c}},\quad\partial_{n}u^{(2)}=-au^{(1)}\quad\textrm{on}~\Gamma_{\mathrm{a}}.$
Integration of the Poisson equation over $\Omega$ yields
$|\Omega|\lambda^{(2)}-\lambda^{(1)}\int\limits_{\Omega}u^{(1)}=\int\limits_{\Gamma_{\mathrm{c}}}c\,u^{(1)}-\int\limits_{\Gamma_{\mathrm{a}}}c\,u^{(1)}.$ (15)
At the same time, multiplying Eq. (14a) by $u^{(1)}$ and integrating over $\Omega$, we deduce with the help of the Green’s formula and boundary conditions:
$\lambda^{(1)}\int\limits_{\Omega}u^{(1)}=-\int\limits_{\Omega}|\nabla u^{(1)}|^{2}+\int\limits_{\Gamma_{\mathrm{c}}}c\,u^{(1)}-\int\limits_{\Gamma_{\mathrm{a}}}a\,u^{(1)}.$ (16)
Subtracting these equations, we obtain
$\lambda^{(2)}=-\frac{1}{|\Omega|}\int\limits_{\Omega}|\nabla u^{(1)}|^{2}<0.$ (17)
We conclude that the second-order correction $\lambda^{(2)}$ to the principal eigenvalue is negative. As a consequence, even if we fix $q_{\mathrm{a}}$ according to Eq. (5) to ensure that $\lambda^{(1)}=0$, the principal eigenvalue $\lambda_{0}$ would be small in amplitude but strictly negative, implying an exponential growth of the mean population size.
Appendix B Solution for the interval
The spectral properties of the Laplace operator on the interval $(0,L)$ are well known (see, e.g., [42, 55, 56]). Searching eigenvalues and eigenfunctions in the form
$u(x)=a\cos(\alpha x)+b\sin(\alpha x),\qquad\lambda=\alpha^{2},$ (18)
one relates the unknown coefficients $a$, $b$ and $\alpha$ by rewriting Robin boundary conditions in Eqs. (2) as
$-b\alpha-q_{\mathrm{c}}a$ $=0\quad(\textrm{at}~x=0),$
$\alpha(-a\sin(\alpha L)+b\cos(\alpha L))$
$+q_{\mathrm{a}}(a\cos(\alpha L)+b\sin(\alpha L))$ $=0\quad(\textrm{at}~x=L).$
Combining these relations yields
$(\alpha^{2}+q_{\mathrm{a}}q_{\mathrm{c}})\frac{\tan(\alpha L)}{\alpha}=q_{\mathrm{a}}-q_{\mathrm{c}},$ (19)
which is equivalent to Eq. (6) from the main text.
In the case $q_{\mathrm{a}}=q_{\mathrm{c}}=q$, we get $(\alpha^{2}+q^{2})\tan(\alpha L)/\alpha=0$ so that $\alpha_{k}L=\pi k$, with $k=1,2,\cdots$, i.e., we retrieve the eigenvalues for an interval with reflecting endpoints. In turn, the principal eigenvalue is given by $\lambda_{0}=\alpha_{0}^{2}=-q^{2}<0$.
Appendix C Critical catalytic rate
Let us consider the confining domain $\Omega=\{\bm{x}\in{\mathbb{R}}^{d}~:~R<|\bm{x}|<L\}$ between two concentric spheres of radii $R$ and $L$ ($d\geq 2$). The inner sphere is the catalytic region with $q_{\mathrm{c}}\geq 0$, whereas the outer sphere is the absorbing region with $q_{\mathrm{a}}\geq 0$. For a given catalytic rate $q_{\mathrm{c}}$, we aim at finding the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ so that the absorption events compensate the branching events and thus give the steady-state population. Setting $\lambda=0$ into Eq. (2), one can search the associated eigenfunction in the radial form as $u(\bm{x})=A+Br^{2-d}$ for $d\geq 3$, or as $u(\bm{x})=A+B\ln r$ for $d=2$, where $r=|\bm{x}|$, with unknown constants $A$ and $B$. The Robin boundary conditions in Eqs. (2) read as
$-\partial_{r}u-q_{\mathrm{c}}u$ $=0\quad\textrm{at}~r=R,$ (20a)
$\partial_{r}u+q_{\mathrm{a}}u$ $=0\quad\textrm{at}~r=L,$ (20b)
allowing us to deduce the relation between $q_{\mathrm{a}}$ and $q_{\mathrm{c}}$.
In the planar case ($d=2$), Eqs. (20) read as
$\frac{B}{q_{\mathrm{c}}R}+A+B\ln R$ $=0,$
$\frac{B}{q_{\mathrm{a}}L}+A+B\ln L$ $=0,$
so that the nontrivial solution is possible if and only if
$\frac{1}{q_{\mathrm{c}}R}+\frac{1}{q_{\mathrm{a}}L}+\ln(L/R)=0,$
from which
$q_{\mathrm{a}}=\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)=q_{\mathrm{c}}\frac{R/L}{1-q_{\mathrm{c}}R\ln(L/R)}\,.$ (21)
We get thus Eq. (7) from the main text, with $q_{\mathrm{c}}^{\rm crit}$ given by Eq. (8).
When $d\geq 3$, Eqs. (20) read as
$\frac{(d-2)B}{q_{\mathrm{c}}R^{d-1}}-\biggl(A+\frac{B}{R^{d-2}}\biggr)$ $=0,$
$-\frac{(d-2)B}{q_{\mathrm{a}}L^{d-1}}+\biggl(A+\frac{B}{L^{d-2}}\biggr)$ $=0,$
so that the nontrivial solution is possible if and only if
$\frac{d-2}{q_{\mathrm{c}}R^{d-1}}-\frac{d-2}{q_{\mathrm{a}}L^{d-1}}+\frac{1}{L^{d-2}}-\frac{1}{R^{d-2}}=0,$
from which
$q_{\mathrm{a}}=\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)=q_{\mathrm{c}}\frac{(R/L)^{d-1}}{1-q_{\mathrm{c}}R(1-(R/L)^{d-2})/(d-2)}\,.$ (22)
We retrieve again Eq. (7) from the main text, with $q_{\mathrm{c}}^{\rm crit}$ given by Eq. (8).
Appendix D Insights from spectral expansions
Additional insights onto the mean population size can be gained from spectral expansions that were obtained in [59, 29] in the case of a single absorbing region with a given reactivity. While an extension of the encounter-based approach to multiple targets was discussed in [60], no spectral expansion is available for this more general case, which includes, in particular, our setting with both catalytic and absorbing regions. In this Section, we propose a partial solution to this problem, which consists in fixing the absorption rate $q_{\mathrm{a}}$ and treating the catalytic rate $q_{\mathrm{c}}$ as a spectral parameter.
Let us consider the Laplace transform of the mean population size:
$\tilde{N}(p|\bm{x})=\int\limits_{0}^{\infty}dt\,e^{-pt}\,N(t|\bm{x}),$ (23)
which, according to Eq. (1), should satisfy the mixed boundary value problem:
$(p-D\Delta)\tilde{N}(p|\bm{x})$ $=1\quad\textrm{in}~\Omega,$ (24a)
$\partial_{n}\tilde{N}(p|\bm{x})$ $=q_{\mathrm{c}}\tilde{N}(p|\bm{x})\quad\textrm{on}~\Gamma_{\mathrm{c}},$ (24b)
$\partial_{n}\tilde{N}(p|\bm{x})+q_{\mathrm{a}}\tilde{N}(p|\bm{x})$ $=0\quad\textrm{on}~\Gamma_{\mathrm{a}},$ (24c)
$\partial_{n}\tilde{N}(p|\bm{x})$ $=0\quad\textrm{on}~\Gamma_{\mathrm{r}}.$ (24d)
One can search the solution of Eqs. (24) as a linear combination of two functions:
$\tilde{N}(p|\bm{x})=\tilde{S}_{0}(p|\bm{x})+\tilde{U}(p|\bm{x}),$ (25)
where $\tilde{S}_{0}(p|\bm{x})$ is the solution of the PDE with the source term in the bulk:
$(p-D\Delta)\tilde{S}_{0}(p|\bm{x})$ $=1\quad\textrm{in}~\Omega,$ (26a)
$\tilde{S}_{0}(p|\bm{x})$ $=0\quad\textrm{on}~\Gamma_{\mathrm{c}},$ (26b)
$\partial_{n}\tilde{S}_{0}(p|\bm{x})+q_{\mathrm{a}}\tilde{S}_{0}(p|\bm{x})$ $=0\quad\textrm{on}~\Gamma_{\mathrm{a}},$ (26c)
$\partial_{n}\tilde{S}_{0}(p|\bm{x})$ $=0\quad\textrm{on}~\Gamma_{\mathrm{r}}.$ (26d)
This solution can be interpreted as the Laplace transform of the survival probability $S_{0}(t|\bm{x})$ for a particle diffusing in $\Omega$ in the presence of the perfectly absorbing region $\Gamma_{\mathrm{c}}$ and partially absorbing region $\Gamma_{\mathrm{a}}$. In other words, this classical quantity does not involve catalytic branching events. In turn, the second function in Eq. (25) is the solution of the PDE with the source term on the boundary:
$(p-D\Delta)\tilde{U}(p|\bm{x})$ $=0\quad\textrm{in}~\Omega,$ (27a)
$\partial_{n}\tilde{U}(p|\bm{x})-q_{\mathrm{c}}\tilde{U}(p|\bm{x})$ $=-\partial_{n}\tilde{S}_{0}(p|\bm{x})\quad\textrm{on}~\Gamma_{\mathrm{c}},$ (27b)
$\partial_{n}\tilde{U}(p|\bm{x})+q_{\mathrm{a}}\tilde{U}(p|\bm{x})$ $=0\quad\textrm{on}~\Gamma_{\mathrm{a}},$ (27c)
$\partial_{n}\tilde{U}(p|\bm{x})$ $=0\quad\textrm{on}~\Gamma_{\mathrm{r}}.$ (27d)
To solve this PDE, we introduce the generalized Steklov problem:
$(p/D-\Delta)V^{(p,q_{\mathrm{a}})}$ $=0\quad\textrm{in}~\Omega,$ (28a)
$\partial_{n}V^{(p,q_{\mathrm{a}})}$ $=\mu^{(p,q_{\mathrm{a}})}V^{(p,q_{\mathrm{a}})}\quad\textrm{on}~\Gamma_{\mathrm{c}},$ (28b)
$\partial_{n}V^{(p,q_{\mathrm{a}})}+q_{\mathrm{a}}V^{(p,q_{\mathrm{a}})}$ $=0\quad\textrm{on}~\Gamma_{\mathrm{a}},$ (28c)
$\partial_{n}V^{(p,q_{\mathrm{a}})}$ $=0\quad\textrm{on}~\Gamma_{\mathrm{r}},$ (28d)
with the spectral parameter $\mu^{(p,q_{\mathrm{a}})}$ instead of $q_{\mathrm{c}}$. In analogy to the conventional Steklov problem (corresponding to $p=0$, $\Gamma_{\mathrm{c}}=\partial\Omega$ and thus $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{r}}=\emptyset$), this spectral problem has a discrete spectrum, i.e., a countable set of eigenvalues $\{\mu_{k}^{(p,q_{\mathrm{a}})}\}$, enumerated by the index $k=0,1,2,\cdots$ in an increasing order, $\mu_{0}^{(p,q_{\mathrm{a}})}\leq\mu_{1}^{(p,q_{\mathrm{a}})}\leq\cdots\nearrow+\infty$; in turn, the restrictions of the associated eigenfunctions $\{V_{k}^{(p,q_{\mathrm{a}})}\}$ onto $\Gamma_{\mathrm{c}}$ form a complete orthonormal basis of $L^{2}(\Gamma_{\mathrm{c}})$ [41]. The last property allows one to expand any function from $L^{2}(\Gamma_{\mathrm{c}})$ on this basis and thus to get the following spectral expansion of $\tilde{U}(p|\bm{x})$:
$\tilde{U}(p|\bm{x})=\sum\limits_{k=1}^{\infty}\frac{V_{k}^{(p,q_{\mathrm{a}})}(\bm{x})}{\mu_{k}^{(p,q_{\mathrm{a}})}-q_{\mathrm{c}}}\int\limits_{\Gamma_{\mathrm{c}}}d\bm{y}\,V_{k}^{(p,q_{\mathrm{a}})}(\bm{y})(-\partial_{n}\tilde{S}_{0}(p|\bm{y})).$ (29)
Since both $\tilde{S}_{0}(p|\bm{x})$ and the eigenpairs $\{\mu_{k}^{(p,q_{\mathrm{a}})},V_{k}^{(p,q_{\mathrm{a}})}\}$ are in general hard to access, this is a very formal representation of the mean population size $N(t|\bm{x})$ in the Laplace domain.
At the same time, this formal representation brings some spectral insights onto the asymptotic behavior of the mean population size. In fact, setting $p/D=-\lambda$ reduces Eqs. (28) to the eigenvalue problem (2) for the Laplace operator. This reflects the duality between the Robin and Steklov spectral problems [41]. In particular, the Laplacian eigenvalues $\lambda_{k}$ satisfying (2) can be searched as real solutions of equations $\mu_{j}^{(-D\lambda,q_{\mathrm{a}})}=q_{\mathrm{c}}$ with $j=0,1,2,\ldots$. This is also consistent with the Laplace transform inversion of the spectral expansion (27) that requires finding the poles $p_{k}=-D\lambda_{k}$ of $\tilde{U}(p|\bm{x})$ that corresponds to zeros of $\mu_{k}^{(p,q_{\mathrm{a}})}-q_{\mathrm{c}}$ (note that the principal eigenvalue $\lambda_{0}$ is related to the principal eigenvalue $\mu_{0}^{(p,q_{\mathrm{a}})}$).
To illustrate this point, let us consider again the spherical shell $\Omega=\{\bm{x}\in{\mathbb{R}}^{3}~:~R<|\bm{x}|<L\}$ between two concentric spheres of radii $R$ and $L$, associated with the catalytic and absorbing regions, respectively (here $\Gamma_{\mathrm{r}}=\emptyset$). The rotational symmetry of this domain allows for the separation of variables to solve Eq. (28). In particular, the eigenfunction $V_{0}^{(p,q_{\mathrm{a}})}$ associated to the principal eigenvalue $\mu_{0}^{(p,q_{\mathrm{a}})}$ can be searched in the form (up to a suitable normalization):
$V_{0}^{(p,q_{\mathrm{a}})}(r)=\biggl[\alpha i^{\prime}_{0}(\alpha L)+q_{\mathrm{a}}i_{0}(\alpha L)\biggr]k_{0}(\alpha r)$
$\qquad-\biggl[\alpha k^{\prime}_{0}(\alpha L)+q_{\mathrm{a}}k_{0}(\alpha L)\biggr]i_{0}(\alpha r)\quad(R\leq r\leq L),$
which respects Eqs. (28a) and (28c). Here $\alpha=\sqrt{p/D}$, $i_{0}(z)=\sinh(z)/z$, $k_{0}(z)=e^{-z}/z$, and prime denotes the derivative with respect to the argument. In turn, the boundary condition (28b) determines the eigenvalue as
$\mu_{0}^{(p,q_{\mathrm{a}})}=\frac{(-\partial_{r}V_{0}^{(p,q_{\mathrm{a}})})_{r=R}}{V_{0}^{(p,q_{\mathrm{a}})}(R)}\,.$ (30)
Figure 3 illustrates the behavior of $\mu_{0}^{(p,q_{\mathrm{a}})}$ as a function of $p$ for three values of $q_{\mathrm{a}}$. Expectedly, all three curves monotonously increase with $p$, and they cannot cross each other. Moreover, $\mu_{0}^{(p,q_{\mathrm{a}})}$ diverges to -$\infty$ as $p$ approaches to some critical value $p_{q_{\mathrm{a}}}$, which depends on $q_{\mathrm{a}}$. Even though $\mu_{0}^{(p,q_{\mathrm{a}})}$ is well defined for $p<p_{q_{\mathrm{a}}}$, we restrict our discussion to $p>p_{q_{\mathrm{a}}}$. The principal Laplacian eigenvalue $\lambda_{0}$ can be obtained as $p_{0}/D$, where $p_{0}$ is the unique solution of the equation $\mu_{0}^{(p,q_{\mathrm{a}})}=q_{\mathrm{c}}$ on $(p_{q_{\mathrm{a}}},\infty)$. Depending on $q_{\mathrm{c}}$ and $q_{\mathrm{a}}$, one may get $p_{0}<0$, $p_{0}=0$ or $p_{0}>0$. If we are interested in the steady-state solution with $\lambda_{0}=0$ (and thus $p_{0}=0$), the eigenvalue $\mu_{0}^{(0,q_{\mathrm{a}})}$ determines the catalytic rate $q_{\mathrm{c}}$, associated to the imposed absorption rate $q_{\mathrm{a}}$. Alternatively, one can formally invert the function $\mu_{0}^{(0,q_{\mathrm{a}})}$ to determine $q_{\mathrm{a}}$ from a given $q_{\mathrm{c}}$. From this figure, it is also clear that if $q_{\mathrm{c}}$ lies above $\mu_{0}^{(0,\infty)}$ (shown by a red filled circle), only a positive solution $p_{0}$ of $\mu_{0}^{(p,q_{\mathrm{a}})}=q_{\mathrm{c}}$ exists, yielding $\lambda_{0}<0$ and thus an exponential growth of the population. In other words, $\mu_{0}^{(0,\infty)}$ determines the critical catalytic rate: $q_{\mathrm{c}}^{\rm crit}=\mu_{0}^{(0,\infty)}$. Moreover, setting $p=0$ and $q_{\mathrm{a}}\to\infty$, one reduces Eqs. (28) to the eigenvalue problem (9) so that $\mu_{0}^{(0,\infty)}$ can be identified with $\mu_{0}$ discussed in the main text.
Figure 3: The principal eigenvalue $\mu_{0}^{(p,q_{\mathrm{a}})}$ from Eq. (30) as a function of $p$ for the spherical shell with $R=0.2$ and $L=1$, $D=1$, and three values of $q_{\mathrm{a}}$ as indicated in the legend. Big red circle indicates the point $\mu_{0}^{(0,\infty)}$ that corresponds to $\mu_{0}$ determining the critical catalytic rate $q_{\mathrm{c}}^{\rm crit}$, see Eqs. (9). Two vertical dash-dotted lines indicate the values $p_{q_{\mathrm{a}}}$, at which $\mu_{0}^{(p,q_{\mathrm{a}})}$ diverges; note that $\mu_{0}^{(p,q_{\mathrm{a}})}$ is not shown for $p<p_{q_{\mathrm{a}}}$.
Appendix E Asymptotic behavior near critical regime
In this Section, we analyze the asymptotic behavior of the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ as the catalytic rate $q_{\mathrm{c}}$ approaches its critical value $q_{\mathrm{c}}^{\rm crit}$. We recall that $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ corresponds to the solution of Eq. (2) with $\lambda=0$ and can be determined as $\sigma_{0}$, where $\sigma_{0}$ is the principal eigenvalue of the following Steklov problem
$\Delta u$ $=0\quad\textrm{in}~\Omega,$ (31a)
$\partial_{n}u$ $=\sigma u\quad\textrm{on}~\Gamma_{\mathrm{a}},$ (31b)
$\partial_{n}u$ $=q_{\mathrm{c}}u\quad\textrm{on}~\Gamma_{\mathrm{c}},\qquad\partial_{n}u=0\quad\textrm{on}~\Gamma_{\mathrm{r}}.$ (31c)
We multiply Eq. (31a) by the eigenfunction $v$ satisfying Eqs. (9), multiply Eq. (9a) by $u$, subtract them, integrate over $\Omega$, use the Green’s formula, and apply the boundary conditions to get
$0$ $=\int\limits_{\Omega}\bigl(u\Delta v-v\Delta u\bigr)=\int\limits_{\partial\Omega}\bigl(u\partial_{n}v-v\partial_{n}u\bigr)$
$=\int\limits_{\Gamma_{\mathrm{a}}}\frac{\partial_{n}u}{\sigma}\partial_{n}v+(\mu-q_{\mathrm{c}})\int\limits_{\Gamma_{\mathrm{c}}}uv.$
Applying this identity to the principal eigenpair $\{\sigma_{0},u_{0}\}$ of the spectral problem (31) and the principal eigenpair $\{\mu_{0},v_{0}\}$ of the spectral problem (9), we identify $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ with $\sigma_{0}$ and $q_{\mathrm{c}}^{\rm crit}$ with $\mu_{0}$. In this case, the above identity reads
$\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)=\frac{1}{q_{\mathrm{c}}^{\rm crit}-q_{\mathrm{c}}}\frac{\int\nolimits_{\Gamma_{\mathrm{a}}}(\partial_{n}u_{0})(\partial_{n}v_{0})}{\int\nolimits_{\Gamma_{\mathrm{c}}}u_{0}v_{0}}\,.$ (32)
This relation is exact and valid for any $0<q_{\mathrm{c}}<q_{\mathrm{c}}^{\rm crit}$. As the eigenfunction $u_{0}$ depends on $q_{\mathrm{c}}$, this representation remains quite formal. However, as the limit $q_{\mathrm{c}}\to q_{\mathrm{c}}^{\rm crit}$ corresponds to $q_{\mathrm{a}}\to\infty$ (and thus $\sigma_{0}\to$-$\infty$), the PDE (31) becomes closer and closer to the PDE (9), so that $u_{0}$ approaches $v_{0}$. As a consequence, the ratio of two integrals in Eq. (32) approaches a finite limit so that
$\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)\approx\frac{1}{q_{\mathrm{c}}^{\rm crit}-q_{\mathrm{c}}}\frac{\int\nolimits_{\Gamma_{\mathrm{a}}}(\partial_{n}v_{0})^{2}}{\int\nolimits_{\Gamma_{\mathrm{c}}}v_{0}^{2}}\qquad(q_{\mathrm{c}}\to q_{\mathrm{c}}^{\rm crit}).$ (33)
We therefore deduced Eq. (10) from the main text.
In a similar way, we can estimate the principal eigenvalue $\lambda_{0}$ when $q_{\mathrm{c}}$ is close to $q_{\mathrm{c}}^{\rm crit}$ (and $q_{\mathrm{a}}=\infty$). Multiplying Eq. (2a) by $v_{0}$, multiplying Eq. (9a) by $u$, subtracting them, integrating over $\Omega$, using the Green’s formula and boundary conditions, we get
$\lambda\int\limits_{\Omega}u\,v_{0}$ $=\int\limits_{\Omega}(u\,\Delta v_{0}-v_{0}\,\Delta u)=\int\limits_{\partial\Omega}(u\,\partial_{n}v_{0}-v_{0}\,\partial_{n}u)$
$=(q_{\mathrm{c}}^{\rm crit}-q_{\mathrm{c}})\int\limits_{\Gamma_{\mathrm{c}}}u\,v_{0},$
where $\mu_{0}$ was again identified with $q_{\mathrm{c}}^{\rm crit}$. This identity is valid for any eigenpair $\{\lambda,u\}$ of the Laplace operator. When $\lambda=\lambda_{0}$ is the principal eigenvalue, the associated eigenfunction $u_{0}$ should be close to $v_{0}$ in the limit $q_{\mathrm{c}}\to q_{\mathrm{c}}^{\rm crit}$, so that
$\lambda_{0}\approx(q_{\mathrm{c}}^{\rm crit}-q_{\mathrm{c}})\frac{\int\nolimits_{\Gamma_{\mathrm{c}}}v_{0}^{2}}{\int\nolimits_{\Omega}v_{0}^{2}}\,.$ (34)
This first-order approximation can be seen as a linear response of the system to a small perturbation of $q_{\mathrm{c}}$ around its critical value $q_{\mathrm{c}}^{\rm crit}$. In particular, if $q_{\mathrm{c}}>q_{\mathrm{c}}^{\rm crit}$, then $\lambda_{0}<0$, and the mean population size grows exponentially.
Appendix F Finite-elements method
In this Section, we briefly describe the numerical technique that we used to prepare Fig. 2. We focus on the steady-state regime and consider two related problems: (i) computation of the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ by solving the Steklov problem (31); (ii) computation of the critical catalytic rate $q_{\mathrm{c}}^{\rm crit}$ by solving the Steklov problem (9).
Both Steklov problems were solved by using a home-built implementation of a finite-element method developed in [43]. All computations are done in Matlab PDEtool, which was also used to triangulate the computational domain $\Omega$. The maximal meshsize $h_{\rm max}$ controls the accuracy of the computation, which was validated on a circular annulus, for which the exact solutions (7, 8) are available.
Appendix G Small-region asymptotic results
In this Section, we employ recent asymptotic results to estimate the critical catalytic rate $q_{\mathrm{c}}^{\rm crit}$ in the case when both catalytic and absorbing regions are small. We recall that $q_{\mathrm{c}}^{\rm crit}$ is given by the principal eigenvalue $\mu_{0}$ of the Steklov problem (9). The asymptotic analysis of this spectral problem in two dimensions was undertaken in [44], which was focused on the case when the smooth boundary $\partial\Omega$ is connected, i.e., both regions $\Gamma_{\mathrm{a}}$ and $\Gamma_{\mathrm{c}}$ are reactive patches on the otherwise reflecting boundary. More precisely, $\Gamma_{\mathrm{c}}$ was assumed to be a single connected arc-like subset, centered around $\bm{x}_{1}\in\partial\Omega$ and of length $2\epsilon_{1}$, whereas $\Gamma_{\mathrm{a}}$ was the union of N1N-1 arc-like subsets, each being connected, centered around $\bm{x}_{j}\in\partial\Omega$ and of length $2\epsilon_{j}$, $j=2,\cdots,N$. All lengths $\epsilon_{j}$ are small as compared to the diameter $2L$ of the domain, $\varepsilon_{j}=\epsilon_{j}/L\ll 1$, whereas the centers $\bm{x}_{i}$ and $\bm{x}_{j}$ are well separated from each other: $|\bm{x}_{i}-\bm{x}_{j}|\sim L$ for all $i\neq j$.
Under these assumptions, the asymptotic behavior of the principal eigenvalue $\mu_{0}$ of the spectral problem (9) was derived in [44] and is given by Eq. (11). The factor $C$ in this equation incorporates all relevant geometric information such as the lengths of all patches, their locations on the boundary $\partial\Omega$, and the shape of the confining domain. It can be written in a compact form as
$C=\frac{2}{\pi}\biggl({\bf{e}}_{1}^{\dagger}{\bf{M}}_{0}^{-1}({\bf{e}}_{1}-\bm{\nu}{\bf{e}}/\bar{\nu})\biggr)^{-1},$ (35)
where ${\bf{M}}_{0}={\bf{I}}+\bigl({\bf{I}}-\bm{\nu}{\bf{e}}{\bf{e}}^{\dagger}/\bar{\nu}\bigr)\bm{\nu}{\bf{G}}$, $\bf{I}$} is the identity matrix, $\nu_{1}=-1/\ln(\varepsilon_{1})$, $\nu_{j}=-1/\ln(\varepsilon_{j}/2)$ for $j=2,3,\ldots,N$, $\bar{\nu}=\nu_{1}+\ldots+\nu_{N}$, and we used the following matrix notations:
$\bm{\nu}=\left(\begin{array}[]{c c c c}\nu_{1}&0&\ldots&0\\ 0&\nu_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&\nu_{N}\\ \end{array}\right),~{\bf{e}}$ $=\left(\begin{array}[]{c}1\\ 1\\ \ldots\\ 1\\ \end{array}\right),~{\bf{e}}_{k}=\left(\begin{array}[]{c}0\\ 1\\ \ldots\\ 0\\ \end{array}\right),$
where $1$ stands on the $k$-th row of the vector ${\bf{e}}_{k}$. The matrix $\bf{G}$} has the elements
${\bf{G}}_{i,j}=\pi G(\bm{x}_{i},\bm{x}_{j})\quad(i\neq j),\qquad{\bf{G}}_{i,i}=\pi R(\bm{x}_{i}),$ (36)
where $G(\bm{x},\bm{y})$ is the surface Neumann Green’s function satisfying for any $\bm{y}\in\partial\Omega$:
$\Delta_{\bm{x}}G(\bm{x},\bm{y})$ $=\frac{1}{|\Omega|}\quad\textrm{in}~\Omega,\qquad\int\limits_{\Omega}G(\bm{x},\bm{y})d\bm{x}=0,$ (37a)
[…]
$G(\bm{x},\bm{y})$ $\underset{\bm{x}\to\bm{y}}{\sim}-\frac{1}{\pi}\ln|\bm{x}-\bm{y}|+R(\bm{y})+o(1),$ (37c)
where $|\Omega|$ is the volume of $\Omega$, and $R(\bm{y})$ is the regular part of $G(\bm{x},\bm{y})$ (i.e., its constant value once the singular logarithmic term is subtracted).
To illustrate this asymptotic result, we take $\Omega$ to be the unit disk ($L=1$), for which
$G(\bm{x},\bm{y})=-\frac{1}{\pi}\ln|\bm{x}-\bm{y}|+\frac{|\bm{x}|^{2}}{4\pi}-\frac{1}{8\pi},\quad R(\bm{y})=\frac{1}{8\pi}\,.$ (38)
If the absorbing region consists of a single patch, $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{a}}^{1}$, the above matrices of size $2\times 2$ can be computed explicitly, yielding Eq. (12), see [44] for details.
As discussed in [44], the above asymptotic results can also be generalized to the case when $\Gamma_{\mathrm{c}}$ and/or $\Gamma_{\mathrm{a}}^{i}$ are located in the bulk of $\Omega$. In fact, one has to set $\nu_{j}=-1/\ln(d_{j})$, where $d_{1}$ is the logarithmic capacity of the catalytic region $\Gamma_{\mathrm{c}}$ and $d_{j}$ is the logarithmic capacity of the absorbing subset $\Gamma_{\mathrm{a}}^{j}$ ($j=2,3,\cdots,N$); in addition, Eq. (36) is replaced by
${\bf{G}}_{i,j}=2\pi G(\bm{x}_{i},\bm{x}_{j})\quad(i\neq j),\quad{\bf{G}}_{i,i}=2\pi R(\bm{x}_{i}),$ (39)
where $G(\bm{x},\bm{y})$ is the bulk Neumann Green’s function (or pseudo-Green’s function) satisfying for any $\bm{y}\in\Omega$:
[…]
$\partial_{n}G$ $=0\quad\textrm{on}~\partial\Omega,\quad\int\limits_{\Omega}G(\bm{x},\bm{y})d\bm{x}=0,$ (40b)
$G(\bm{x},\bm{y})$ $\underset{\bm{x}\to\bm{y}}{\sim}-\frac{1}{2\pi}\ln|\bm{x}-\bm{y}|+R(\bm{y})+o(1),$ (40c)
and $R(\bm{y})$ is its regular part. Finally, the constant $C_{0}$ has to be recomputed from the asymptotic behavior of the reactive capacitance ${\mathcal{C}}(\mu)$ (see Appendix C in [44] for details).
While a systematic analysis of this problem is beyond the scope of this paper, we provide an extension of Eq. (12) to the particular case when the catalytic region $\Gamma_{\mathrm{c}}$ is the circle of radius $\epsilon_{1}$, located at point $\bm{x}_{1}\in\Omega$, whereas the absorbing region $\Gamma_{\mathrm{a}}$ is the union of N1N-1 circles of radii $\epsilon_{j}$, located at points $\bm{x}_{j}\in\Omega$. As previously, we assume that all $N$ circles are small ($\epsilon_{j}\ll L$) and well-separated from each other and from the reflecting outer boundary $\Gamma_{\mathrm{r}}$ of $\Omega$ (i.e., $|\bm{x}_{i}-\bm{x}_{j}|\sim L$ and $|\bm{x}_{i}-\Gamma_{\mathrm{r}}|\sim L$), where $2L$ is the diameter of the domain $\Omega$. In this case, $d_{j}=\epsilon_{j}$, ${\mathcal{C}}(\mu)=1/\mu$ so that Eq. (5.13) from [44] becomes
$\frac{1}{\mu_{0}\epsilon_{1}}\approx\frac{1}{\nu_{1}}\biggl({\bf{e}}_{1}^{\dagger}{\bf{M}}_{0}^{-1}({\bf{e}}_{1}-\bm{\nu}{\bf{e}}/\bar{\nu})\biggr).$ (41)
As a consequence, Eq. (12) remains valid, with
$C=\biggl({\bf{e}}_{1}^{\dagger}{\bf{M}}_{0}^{-1}({\bf{e}}_{1}-\bm{\nu}{\bf{e}}/\bar{\nu})\biggr)^{-1},\quad C_{0}=0.$ (42)
In the case when $\Omega$ is the unit disk ($L=1$), the bulk Neumann Green’s function was given in [57]:
$G(\bm{x},\bm{y})$ $=\frac{1}{2\pi}\biggl(-\ln|\bm{x}-\bm{y}|-\ln\biggl|\bm{x}|\bm{y}|-\bm{y}/|\bm{y}|\biggr|$
$+\frac{1}{2}(|\bm{x}|^{2}+|\bm{y}|^{2})-\frac{3}{4}\biggr),$ (43a)
$R(\bm{y})$ $=\frac{1}{2\pi}\biggl(-\ln(1-|\bm{y}|^{2})+|\bm{y}|^{2}-\frac{3}{4}\biggr).$ (43b)
For instance, in the case of two circular regions (one catalytic and one absorbing), we get (see Eq. (6.11) from [44])
$q_{\mathrm{c}}^{\rm crit}\simeq\frac{1}{\epsilon_{1}\bigl[-\ln(\epsilon_{1}\epsilon_{2})+\delta G(\bm{x}_{1},\bm{x}_{2})\bigr]}\,$, (44)
where
$\delta G(\bm{x}_{1},\bm{x}_{2})=2\pi[R(\bm{x}_{1})+R(\bm{x}_{2})-2G(\bm{x}_{1},\bm{x}_{2})]$
$=\ln\biggl(\frac{|\bm{x}_{1}-\bm{x}_{2}|^{2}\,\bigl|\bm{x}_{1}|\bm{x}_{2}|-\bm{x}_{2}/|\bm{x}_{2}|\bigr|^{2}}{(1-|\bm{x}_{1}|^{2})(1-|\bm{x}_{2}|^{2})}\biggr).$ (45)
To illustrate the utility of these asymptotic results, we estimate the critical catalytic rate $q_{\mathrm{c}}^{\rm crit}$ from Eqs. (12, 42) for two configurations shown in Fig. 2(a,b). Substituting the centers $\bm{x}_{j}$ of ten circles into Eqs. (39, 43), we first evaluate the matrix $\bf{G}$}. Setting $\epsilon_{1}=R=0.1$ and $\epsilon_{j}=R_{\mathrm{a}}=0.05$ ($j=2,\cdots,10$) for the first configuration, we compute the matrices $\bm{\nu}$ and ${\bf{M}}_{0}$, from which we get the coefficient $C$ from Eq. (42). Its substitution into Eq. (12) yields $q_{\mathrm{c}}^{\rm crit}\approx 6.57$, which is remarkably close to the numerical value $q_{\mathrm{c}}^{\rm crit}\approx 6.61$ obtained by solving the spectral problem (9) by a FEM (see Appendix F). Even though the circles are not too small, the relative error of the asymptotic value is below $1\%$. For the second configuration with even larger circles ($R_{\mathrm{a}}=0.1$), the asymptotic value is $7.31$, which differs from the benchmark value $7.62$ of the FEM by $4\%$. A further analysis of critical catalytic rates in complex environments with multiple small absorbing regions presents an interesting perspective. An extension of the asymptotic results to the three-dimensional case is also possible (see [45]).
Appendix H Solution for the exterior of a ball
It is instructive to consider diffusion in the exterior of a ball of radius $R$: $\Omega=\{\bm{x}\in{\mathbb{R}}^{3}~:~|\bm{x}|>R\}$. If the spherical boundary is partially absorbing with the rate $q_{\mathrm{a}}$, the solution of Eq. (1) is the survival probability of a particle, whose explicit solution in the context of chemical physics was first discussed in [58]:
$S(t|\bm{x})=1-\frac{R/r}{1+1/(q_{\mathrm{a}}R)}e^{-(r-R)^{2}/(4Dt)}\times$ (46)
$\biggl\{\mathrm{erfcx}\biggl(\frac{r-R}{\sqrt{4Dt}}\biggr)-\mathrm{erfcx}\biggl(\frac{r-R}{\sqrt{4Dt}}+\sqrt{Dt}(q_{\mathrm{a}}+1/R)\biggr)\biggr\},$
where $r=|\bm{x}|$, and $\mathrm{erfcx}(z)=e^{z^{2}}\mathrm{erfc}(z)$ is the scaled complementary error function (note that Eq. (29) from [58] contains a misprint: the sign in front of the last term should be negative, as in Eq. (46)). The expression (46) remains valid for the boundary-catalytic branching on the spherical surface if the absorption rate $q_{\mathrm{a}}$ is replaced by the negative catalytic rate $q_{\mathrm{c}}$:
$N(t|\bm{x})=1-\frac{R/r}{1-1/(q_{\mathrm{c}}R)}e^{-(r-R)^{2}/(4Dt)}\times$ (47)
$\biggl\{\mathrm{erfcx}\biggl(\frac{r-R}{\sqrt{4Dt}}\biggr)-\mathrm{erfcx}\biggl(\frac{r-R}{\sqrt{4Dt}}+\sqrt{Dt}(-q_{\mathrm{c}}+1/R)\biggr)\biggr\}.$
To inspect its long-time limit, we use that $\mathrm{erfcx}(z)\approx 1$ as $z\to 0$, $\mathrm{erfcx}(z)\approx 1/(z\sqrt{\pi})$ as $z\to\infty$, and $\mathrm{erfcx}(z)\simeq 2e^{z^{2}}$ as $z\to$-$\infty$. Using $q_{\mathrm{c}}^{\rm crit}=1/R$ from Eq. (8), we distinguish three regimes.
(i) When $q_{\mathrm{c}}<1/R$, the mean population size $N(t|\bm{x})$ approaches the steady-state limit
$N(\infty|\bm{x})=1-\frac{R/r}{1-1/(q_{\mathrm{c}}R)}\,.$ (48)
As $q_{\mathrm{c}}$ is positive, the branching events lead to the mean population size above $1$. Moreover, as $q_{\mathrm{c}}\to 1/R$, the limiting value $N(\infty|\bm{x})$ diverges, thus indicating that the branching events become more and more proliferative.
(ii) When $q_{\mathrm{c}}>1/R$, the last term in Eq. (47) leads to the exponential growth:
$N(t|\bm{x})\simeq\frac{2R/r}{1-1/(q_{\mathrm{c}}R)}\,e^{Dt(q_{\mathrm{c}}-1/R)^{2}}\quad(t\to\infty),$ (49)
in agreement with our general analysis; in particular, the isolated eigenvalue of the Laplace operator is $\lambda_{0}=-(q_{\mathrm{c}}-1/R)^{2}$.
(iii) Finally, at the critical catalytic rate, $q_{\mathrm{c}}=1/R$, one can evaluate the limit $q_{\mathrm{c}}\to 1/R$ in Eq. (47) to get
$N(t|\bm{x})$ $=1+\frac{2\sqrt{Dt}}{\sqrt{\pi}\,r}e^{-(r-R)^{2}/(4Dt)}$
$-(1-R/r)\mathrm{erfc}\biggl(\frac{r-R}{\sqrt{4Dt}}\biggr).$ (50)
In the long-time limit, one has $N(t|\bm{x})\simeq 2\sqrt{Dt}/(r\sqrt{\pi})$ in the leading order.